### v2

# Effective Online Exam Proctoring by Combining Lightweight Face Detection and Deep Recognition

Abstract
Online exams, conducted via video conferencing platforms such as Zoom, have become popular in educational institutions since COVID-19. While convenient, ensuring the integrity and security of online exams remains challenging, as traditional invigilation struggles to effectively monitor multiple student video feeds in real time. In this paper, we present iExam, an effective online exam proctoring and analysis system that combines lightweight face detection and deep recognition. iExam employs real-time face detection to assist invigilators in continuously monitoring student presence, and leverages deep face recognition for post-exam video analysis to identify abnormal behaviors—including face disappearance, face rotation, and identity substitution. To realize this system, we address three core challenges: (i) designing a lightweight approach to efficiently capture and analyze exam video streams in real time; (ii) developing an enhanced OCR method to automatically extract student identities from dynamically positioned Zoom name tags, enabling reliable ground truth labeling without manual intervention; and (iii) optimizing the training and inference pipeline to significantly reduce resource and time requirements on ordinary teacher devices. Extensive experiments demonstrate that iExam achieves 90.4% accuracy for real-time face detection and 98.4% accuracy for post-exam face recognition, while maintaining low overhead. These results show that iExam can substantially enhance the automation and reliability of online exam proctoring in practice.
I Introduction
The rapid adoption of online and hybrid education, especially in the wake of the COVID-19 pandemic, has fundamentally changed the landscape of teaching and assessment [chakraborty2021opinion, bozkurt2022impact, pokhrel2021literature]. Synchronous online exams delivered over video conferencing platforms such as Zoom and Microsoft Teams have become commonplace in schools and universities worldwide [linden2021zoom, microsoft_teams_exam]. While online exams offer great flexibility and accessibility, ensuring their integrity and fairness remains a critical challenge [butler2020systematic, standard_online_exam_cheats, swissinfo_online_exam_concerns, holden2021academic, langenfeld2020internet].
A central difficulty lies in the effective monitoring—or invigilation—of remote exams. Unlike traditional classroom settings, where instructors can physically oversee all students’ behavior, online environments require proctors to supervise dozens of simultaneous video streams, typically presented in a crowded grid of small windows [lee2022online]. Manual monitoring is inherently limited by human attention and is prone to oversight, bias and fatigue, especially during large-scale or extended assessments [belzak2024measuring].
Existing approaches to online proctoring fall broadly into two categories. Commercial automated proctoring systems leverage computer vision and AI to analyze student behavior, but they often depend on intrusive data collection, centralized cloud processing, or require complex installation and configuration [proctoru_official_website, balash2021examining, balash2023educators]. These factors raise privacy concerns and limit adoption in privacy-sensitive or resource-constrained environments. In contrast, manual video monitoring remains widely used but is inherently unscalable and susceptible to subjective errors [pongchaikul2024remote].
To address these limitations, we present iExam, a novel and practical online exam proctoring system that combines lightweight face detection for real-time monitoring with deep learning-based face recognition for post-exam analysis. The core insight behind iExam is to decouple instant invigilation from in-depth integrity checks: lightweight detection runs efficiently in the browser for instant feedback, while deep recognition is deferred to offline processing for maximum accuracy and flexibility.
During the exam, iExam employs modern Web APIs to capture the Zoom gallery view and perform face detection entirely on the client side [getdisplaymedia_api, opencvjs], thereby preserving privacy and enabling scalable real-time performance without transmitting sensitive images to external servers. After the exam, iExam’s post-processing module leverages state-of-the-art Deep Convolutional Neural Networks (DCNNs) [krizhevsky2012imagenet, he2016deep] to recognize faces and flag anomalies, such as face disappearance, rotation, or identity substitution, from recorded video streams. To automate the collection of ground truth labels for training, we further introduce an enhanced Optical Character Recognition (OCR) technique that reliably extracts student names from dynamic on-screen labels [tesseract_github, yaqub2023proctoring].
To ensure that iExam remains accessible and practical for real-world deployment, for example on standard teacher laptops, which often involves limited computational resources, we incorporate a series of efficiency-oriented optimizations that significantly reduce training and computational overhead, without sacrificing accuracy. Key techniques include sparse OCR labeling, which minimizes redundant text extraction operations; multi-threaded processing, enabling smooth performance even on modest hardware; and automated model selection, which dynamically balances speed and performance based on available resources. These optimizations allow robust face detection algorithms and deep convolutional neural networks to be executed on lightweight hardware, eliminating the need for dedicated GPU servers or costly cloud infrastructure. By keeping both detection and recognition pipelines locally deployable, iExam offers a privacy-preserving and resource-efficient solution for online exam proctoring.
We validate iExam through a series of comprehensive experiments on real-world exam datasets collected from university courses. Results show that our system consistently achieves high accuracy in both real-time face detection (over 90%) and post-exam recognition (about 98%), with acceptable runtime performance and scalability. Moreover, instructors benefit from automated visual reports and anomaly summaries that not only reduce manual monitoring workload but also provide auditable exam integrity checks, improving trust and transparency in online assessments. These results highlight that iExam is a deployable and effective solution bridging the gap between lightweight real-time proctoring and rigorous post-exam analysis.
Contributions. In sum, our contributions are summarized as follows:
  • We design and implement iExam, a two-stage online exam proctoring system that decouples lightweight, privacy-preserving real-time face detection from deep learning-based post-exam face recognition and anomaly analysis.
  • We develop an automated ground truth labeling pipeline using enhanced OCR techniques, enabling reliable extraction of student identities from dynamic Zoom window labels without manual annotation.
  • We introduce a set of resource-efficient training and inference optimizations, including sparse OCR labeling, batch and multi-threaded processing, and model selection strategies, allowing our system to run effectively on standard teacher hardware.
  • We conduct comprehensive experiments with real-world online exam videos, demonstrating that iExam achieves high accuracy (over 90% in real-time detection and 98% in post-exam recognition) and delivers interpretable, actionable reports to instructors for scalable and trustworthy online exam supervision.
II Background
In this section, we introduce the typical online exam scenario using Zoom as the platform, and review mainstream face detection and recognition algorithms.
II-A Online Exams in the “Zoom” Setting
Zoom is one of the most widely adopted platforms for online education and remote examinations [linden2021zoom], providing real-time video conferencing and recording functionalities. In its “gallery view” mode, Zoom displays up to 25 participants per screen (5 rows $\times$ 5 columns) by default, and up to 49 participants (7 $\times$ 7) with modified settings. When the number of exam-takers exceeds the single-screen capacity, additional assistants are required to record multiple screens in parallel. Each session can be recorded and saved as an “mp4” video file, preserving all participants’ facial information, which serves as the input for our iExam system.
A unique feature of Zoom is the personalized label—each participant’s name is shown in the lower left corner of their video window (see Fig. 1), which provides a potential anchor for automated identity extraction. While Zoom enables invigilators to monitor students’ surroundings, it lacks built-in algorithms to detect face disappearance, abnormal behavior, or potential identity substitution, necessitating the development of automated proctoring solutions.
Figure 1: Zoom gallery view with participant video feeds and name labels (masked for anonymity), serving as the input for automated face detection and identity recognition in iExam.
II-B Mainstream Face Classification Approaches
Over the years, researchers have proposed a variety of algorithms for face classification, which can be broadly categorized into feature-based and image-based methods [kumar2019face].
Feature-based approaches exploit properties such as gray-scale values, color, or edges for face detection. In the image-based category, linear subspace techniques, such as Eigenface [turk1991eigenfaces, turk1991face], Principal Component Analysis (PCA) [sirovich1987low, ramadhani2017human], and Linear Discriminant Analysis (LDA) [chan2008face], have shown promise for face recognition by projecting facial images into lower-dimensional subspaces.
However, recent advances in deep learning, especially convolutional neural networks (CNNs) [howard2017mobilenets, huang2017densely, SqueezeNet, szegedy2015going], have significantly improved both the accuracy and robustness of face detection and recognition. Deep Convolutional Neural Networks (DCNNs) [lecun1998gradient, krizhevsky2012imagenet, he2016deep] represent a more advanced class of models that extend traditional CNN architectures to achieve higher accuracy and robustness. Some models [deng2020retinaface, zhang2016joint, wang2023efficientface], pre-trained on large-scale face datasets, achieve lower detection error rates and faster inference compared to traditional algorithms, making them especially suitable for automated online exam proctoring.
III System Overview
In this section, we present a high-level overview of iExam, highlighting its core objectives, architectural rationale, and system workflow. Fig. 2 illustrates the overall architecture.
Figure 2: The overall workflow of iExam, combining lightweight real-time proctoring and deep post-exam analysis.
The key design insight behind iExam is to combine lightweight face detection for real-time monitoring with deep face recognition for accurate, retrospective analysis, achieving both efficiency and robustness for large-scale online exam proctoring. This two-phase architecture is motivated by the distinct requirements at different stages of an online exam:
  • During the exam, real-time proctoring requires rapid feedback on student presence and attention. Here, a lightweight face detection module runs efficiently in the browser, enabling invigilators to promptly identify missing or inattentive students, while minimizing system latency and resource consumption.
  • After the exam, post-exam analysis demands greater accuracy and richer behavioral assessment. Full exam recordings are processed offline using deep learning-based face recognition, which can robustly verify student identities and detect sophisticated anomalies such as face disappearance, rotation, or substitution.
Based on this insight, iExam is composed of two tightly-coupled components:
1) Real-Time Proctoring: iExam provides a web-based platform that captures the Zoom gallery view and applies fast face detection to each video feed in real time. Its primary purpose is to deliver immediate warnings to invigilators — for example, alerting them when a student’s face is absent or turned away from the screen, so that potential issues can be addressed promptly on the spot while efficiently monitoring multiple students simultaneously.
2) Post-Exam Analysis: After the exam, iExam extracts face data from a short training video sampled from the tested video, trains a deep recognition model for each student, and then analyzes the entire recorded session. Unlike real-time proctoring, this component focuses on more in-depth detection of sophisticated anomalies, such as identity substitution. It then generates detailed summary reports, enabling retrospective review and further investigation.
This two-phase design allows iExam to support both real-time invigilation and retrospective integrity verification, providing a robust solution for scalable online exam proctoring.
IV Lightweight Real-time Video Monitoring based on Face Detection
To enable remote, low-cost, and low-latency exam monitoring with minimal configuration, iExam offers a fully online real-time proctoring component optimized for fast and efficient face detection at the edge. This section details the design motivations, key challenges, and implementation details of the lightweight monitoring platform.
Design Challenges. Developing such a system involves three primary challenges: (i) Screen capture: Zoom does not provide a public API for direct video stream access, requiring alternative methods to capture exam sessions; (ii) Detection algorithm selection: achieving a balance between detection accuracy and the lightweight computation necessary for real-time operation; (iii) Computation placement: ensuring that face detection can be performed efficiently and securely, ideally on the client side, to address privacy and network latency concerns.
To address screen capture, we leverage modern browser Web APIs such as getDisplayMedia, allowing users to select and authorize capture of specific application windows (e.g., the Zoom gallery view) as a MediaStream. Our web-based platform thus obtains the live video feed and renders it for in-browser analysis.
For face detection, we offer two JavaScript implementations via OpenCV.js [opencvjs]: (1) the classic Haar-cascade method for rapid, lightweight detection, and (2) a modern DCNN detector for higher accuracy. Both methods can be dynamically loaded and run in the user’s browser, enabling flexible trade-offs between speed and performance depending on hardware capabilities.
A key advantage of iExam is its privacy-preserving, fully client-side design. All face detection is performed locally within the browser—no video frames or facial features are ever transmitted externally. This edge-based architecture eliminates privacy risks inherent to cloud processing and enables seamless, scalable deployment in real-world educational settings.
Figure 3: Web-based proctoring interface: the invigilator selects the Zoom gallery window for monitoring using the browser’s screen capture feature.
Figure 4: Real-time face detection result: detected face regions are overlaid on the mirrored Zoom feed in the browser, providing instant feedback for invigilators.
User Workflow. During an online exam, invigilators open the web-based iExam platform and select the Zoom window to monitor (Fig. 3). The system mirrors the chosen window and overlays real-time face detection results, displaying both the original and processed video streams (Fig. 4). The face detection model is loaded on demand, and all computation occurs locally, ensuring fast and seamless operation without external dependencies.
This lightweight, privacy-preserving monitoring solution empowers teachers to efficiently supervise online exams, detecting student presence and basic anomalies with minimal setup and no special hardware or server infrastructure.
V Comprehensive Post-exam Video Analysis based on Face Recognition
The post-exam analysis module of iExam is designed to automatically detect student absence, face disappearance, and potential identity substitution in recorded exam videos, providing invigilators with concrete and actionable evidence of irregularities. iExam can run cross-platform (e.g., Windows, macOS), and results are delivered through summary reports, including visualizations such as histograms showing face presence over time for each participant.
The workflow of this component consists of three main stages: (1) face dataset construction, (2) classifier training, and (3) face monitoring and feedback generation. To initiate the process, the proctor provides a short sample video (e.g., five minutes), which serves as the training set for individualized face model building.
To accommodate resource constraints typical on teachers’ personal computers, all detected face crops are saved locally as image files and organized by student identity. An adaptive DCNN, such as ResNet [he2016deep], is then trained to extract robust facial features and enable high-accuracy recognition. Once trained, this model processes the entire exam video, associating detected faces with predicted student names and flagging intervals when a participant is absent or unrecognized.
In this task, two primary challenges must be addressed:
  • Automated Labeling: Each face image in the training dataset must be reliably labeled. However, the dynamic layout of Zoom gallery mode means student locations constantly change, making static position-based labeling infeasible.
  • Efficient Model Training and Inference: The system must be able to rapidly train and test deep learning models on commodity teacher devices, balancing recognition accuracy with limited computational resources.
To tackle these, iExam leverages computer vision and OCR (optical character recognition) techniques, paired with multi-threaded processing to accelerate dataset construction and model inference. Fig. 2 illustrates the overall workflow.
V-A Automatic Ground Truth Collection
A major challenge is to automatically obtain labeled face images from the training video, despite students’ positions in the Zoom gallery being dynamic. Our solution is as follows:
Frame Extraction and Face Detection. As the video is processed frame by frame, we utilize a pre-trained DCNN-based face detector, namely Single Shot-Multibox Detector (SSD) [liu2016ssd] with ResNet-10 backbone [he2016deep], resizing extracted face regions to $300\times 300$ pixels for model compatibility. Alongside face extraction, the (x, y, width, height) of each face is recorded to facilitate subsequent cropping and analysis.
The Limitation of Static Position Labeling. The Zoom gallery mode arranges participants in a $5\times 5$ grid, but their order is not fixed. As students join, leave, or lose connection, the positions are shuffled—newcomers are usually placed at the lower right, while others move to fill vacant spots. Thus, it is impossible to rely on a static mapping between cell position and student identity.
Enhancing OCR for Label Extraction. Zoom displays each participant’s name in the lower left of their cell, and this label moves together with the video cell as its position changes. We leverage this observation by using OCR to automatically extract names from these label regions. This provides a robust and dynamic way to assign ground truth labels to cropped face images.
Figure 5: Raw (left) and improved (right) OCR accuracy after thresholding for a participant name.
However, accurate OCR is not trivial: background lighting and low contrast can reduce text recognition accuracy (see left of Fig. 5). To address this, we apply adaptive thresholding to enhance contrast between characters and the background, as illustrated in the right of Fig. 5. We further enlarge and enhance cropped label regions, running OCR at multiple resolutions to maximize recognition accuracy.
Post-processing and Dataset Cleaning. After all candidate face images and labels are collected, the dataset is automatically cleaned. Subfolders (i.e., student classes) with too few images (e.g., less than 100) are treated as noise, since repeated OCR errors for the same person can create multiple small or incorrect subfolders. We also merge folders with similar names (using substring or fuzzy matching with the official roster) to ensure robust label assignment, and prune any folders not corresponding to actual participants. If the video is too short and results in too few ground truth samples, the system prompts the user to provide a longer video for effective training.
V-B Effective and Efficient Face Training and Recognition
The second challenge is to efficiently train and deploy robust face recognition models on standard teacher hardware, while ensuring accuracy and fast turnaround. To address this, we systematically evaluated mainstream face detection and recognition approaches, as well as implemented a series of practical optimizations.
Algorithm Selection and Empirical Comparison. We began by benchmarking several popular face detection methods, including Haar-cascade [viola2001rapid], Dlib [dlib_library], Multi-task Cascaded Convolutional Networks (MTCNNs) [zhang2016joint], and Deep Convolutional Neural Networks (DCNNs). As shown in Fig. 6, the DCNN-based detector achieves both the highest detection count (129,359) and the lowest time cost (196 seconds) among all methods evaluated. MTCNNs and Haar-cascade yield slightly fewer detections, while Dlib records the lowest detection amount. In terms of efficiency, MTCNNs is by far the slowest (3,167 seconds), followed by Dlib (1,683 seconds) and Haar-cascade (960 seconds). These results demonstrate that DCNNs not only offers superior detection accuracy but also delivers significantly faster processing, making it the most suitable choice for our real-time and large-scale post-exam analysis requirements. Based on these findings, we adopted DCNN-based detection as the foundation for subsequent recognition tasks.
Figure 6: Comparison of mainstream face detection algorithms based on a five-minute video.
Model Training and Selection. We evaluated a range of state-of-the-art Convolutional Neural Networks(CNN) architectures, including AlexNet [krizhevsky2012imagenet], GoogleNet [szegedy2015going], various ResNet variants (ResNet18, ResNet50, ResNet152) [he2016deep], as well as SqueezeNet [SqueezeNet], MobileNet [howard2017mobilenets], and DenseNet121 [huang2017densely], using cross-entropy loss as the training objective. To mitigate overfitting arising from the temporal coherence of consecutive video frames, we performed data augmentation such as random flipping, rotation, and cropping during training. This increases the diversity of training samples and enhances the model’s generalization ability.
Figure 7: Training loss curves for different DCNN architectures. All models converge rapidly, typically within 10 epochs.
As shown in Fig. 7, all tested DCNN architectures exhibit rapid convergence, with training loss values dropping sharply within the first few epochs and stabilizing at a low value after approximately 10 epochs. This indicates that the networks can effectively fit the dataset in a short training period, enabling us to employ early stopping strategies to save computation time without sacrificing accuracy. In practice, we found that further increasing the number of epochs yielded diminishing returns in loss reduction. These results demonstrate that mainstream DCNN architectures are well-suited to our dataset and application scenario, and our data augmentation strategies successfully prevent overfitting even when large numbers of highly similar images are present. We ultimately selected the best-performing model, i.e., ResNet50, based on validation accuracy and inference speed in our experiments.
Pipeline Optimizations. To ensure that our face recognition system is both practical and scalable for real-world deployment, we designed a series of optimizations throughout the training and inference pipeline. These enhancements are aimed at minimizing computation and memory overhead, reducing response latency, and making full use of available hardware resources, all while maintaining high recognition accuracy.
  • Sparse OCR Labeling: To accelerate dataset construction, we perform OCR labeling only once per second for each cell, instead of every frame (as videos typically exceed 20 FPS). The captured name is used as the tag for the entire second, significantly reducing the number of OCR operations and the dataset size without sacrificing accuracy.
  • Batch Processing and Multi-threading: Prior to training, all images are shuffled and loaded in batches. We also increase the number of worker threads for the data loader, and enable multi-threaded operations in all stages to fully exploit CPU parallelism.
  • Testing Efficiency: During inference, predictions are buffered for each position every second, rather than for every frame, which greatly reduces redundant computation. A softmax probability threshold is applied so that only predictions above this value are considered valid, reducing false positives.
Robustness to Recognition Failure. If a student’s face is not recognized more than 10 times within a default 30-second interval, the system flags this as a period of disappearance in the report. This design guards against both incidental recognition failures and more severe cases such as proxy attendance or head-turning away from the screen.
Overall, our systematic evaluation and multi-level optimizations allow iExam to deliver effective post-exam face recognition, enabling efficient deployment on educational hardware.
VI Evaluation
VI-A The Performance and Accuracy of Face Detection Algorithms in Real-time Exam Monitoring
To assess the practical effectiveness of iExam’s real-time proctoring module, we conducted a comprehensive evaluation of two widely-used face detection algorithms, i,e, Haar-cascade and DCNNs, under realistic online exam conditions.
Experimental Setup. We collected data from an actual online exam session, resulting in a two-hour video featuring the Zoom gallery view, as depicted in Fig. 3 and 4 (see Sec. IV). For this experiment, we implemented two versions of iExam’s real-time monitoring system: one using the traditional Haar-cascade face detector, and the other employing a modern DCNN-based detector. Apart from the face detection module, all other components and settings, including browser environment and platform configurations, remained identical.
(a) Runtime performance
(b) Detection accuracy
Figure 8: Comparison between DCNNs and Haar-cascade face detection algorithms. Face detection times refer to the total number of face recognitions performed by the method over the entire two-hour video.
Runtime Performance. We first measured the detection efficiency of each face detection algorithm, focusing on the number of successful detections and the corresponding processing time. As shown in Fig. 8(a), Haar-cascade achieves about twice the detection rate of DCNN on the same video stream, primarily due to its simpler, feature-based approach and lower computational requirements. However, this higher speed comes at the expense of accuracy and robustness, as the Haar-cascade method is more prone to missing faces under challenging conditions (e.g., large head rotations, partial occlusions).
In contrast, while the DCNN-based detector processes frames at a lower rate—requiring roughly twice as long to analyze each image—it consistently provides more reliable detection results, especially in cases where faces are tilted, partially covered, or presented at non-frontal angles. This robustness is attributed to the learning-based nature of DCNNs, which enables them to generalize beyond the fixed feature templates used in traditional methods.
Detection Accuracy. To quantify detection accuracy, we conducted a manual evaluation by randomly sampling 50 frames from the two-hour video and annotating the ground truth for each student’s face. Fig. 8(b) illustrates the results, showing that the DCNN detector achieves approximately 90% detection accuracy, significantly outperforming the Haar-cascade method. Notably, the DCNN approach is better able to adapt to a variety of challenging viewing conditions, such as faces appearing close to the camera, rotated, or partially obscured. In contrast, Haar-cascade often fails to detect faces that deviate from standard frontal poses or are occluded, and sometimes misidentifies background objects as faces.
These results highlight a fundamental trade-off in real-time video monitoring: while Haar-cascade is faster, it lacks the robustness and accuracy necessary for reliable online exam proctoring. The DCNN-based detector, despite its relatively higher computational cost, provides the level of precision and adaptability required to maintain exam integrity in practical deployments. For this reason, iExam adopts the DCNN-based approach as the default for real-time invigilation.
VI-B The Performance and Accuracy of Face Recognition Models in Post-exam Video Analysis
In this section, we present a thorough evaluation of iExam’s post-exam analysis component, focusing on the training efficiency and recognition accuracy of eight mainstream DCNN-based face recognition models.
Experimental Setup. We first partition the available face image dataset into three disjoint subsets: 70% for training, 20% for validation, and the remaining 10% for testing. All file indices are randomly shuffled prior to splitting to avoid selection bias. To further verify the robustness of our models, we also evaluate on an independent 90-minute exam video not used during training. The distribution of the training and validation samples is illustrated in Fig. 9. All training experiments were performed on a laptop equipped with an NVIDIA GTX 1650Ti GPU (4GB VRAM), simulating a practical deployment environment for educational institutions.
Figure 9: The distribution of training and validation datasets. Real names were masked and replaced with standardized labels to preserve privacy.
Training Performance. We systematically compare the training time and recognition accuracy of eight popular DCNN architectures, including AlexNet, GoogleNet, ResNet18, ResNet50, ResNet152, SqueezeNet, MobileNet, and DenseNet121. As shown in Fig. 10, GoogleNet stands out as the fastest model, requiring only 244 minutes to complete 10 epochs of training on our dataset. AlexNet and ResNet18 also demonstrate relatively fast training speeds, completing in 293 and 303 minutes, respectively. In contrast, deeper architectures such as ResNet152 and SqueezeNet require substantially more time—650 and 555 minutes, respectively—reflecting the increased computational complexity of larger models. MobileNet and DenseNet121 fall in the mid-to-upper range, with training times of 469 and 445 minutes, respectively.
Figure 10: Training time cost and recognition accuracy of various DCNN models.
Recognition Accuracy. Despite the differences in training efficiency, we observe notable variation in the face recognition accuracy achieved by each model, as summarized in the right-hand section of Fig. 10. ResNet50 achieves the highest accuracy at 98.41%, establishing a clear lead over other models. ResNet152 and GoogleNet also perform well, reaching 93.02% and 88.28% accuracy, respectively. AlexNet and MobileNet deliver moderate results, with accuracies of 77.26% and 78.33%. SqueezeNet lags behind with the lowest accuracy at 69.07%, likely due to its lightweight architecture and reduced parameter count. DenseNet121 and ResNet18 achieve intermediate performance (86.17% and 83.82%, respectively).
This comprehensive evaluation highlights an important trade-off between training time and recognition accuracy: while models such as GoogleNet and AlexNet train rapidly, their accuracy lags behind deeper architectures such as ResNet50. Conversely, extremely deep or parameter-heavy networks (e.g., ResNet152, SqueezeNet) incur longer training times without necessarily delivering superior accuracy. Balancing recognition accuracy and training efficiency is crucial for deployment in real-world online exam scenarios. Based on our experimental results, ResNet50 emerges as the optimal choice for the post-exam analysis component of iExam. Considering that this component is executed offline, we prioritize accuracy over training time. ResNet50 achieves state-of-the-art accuracy (98.41%) while maintaining manageable training time (405 minutes). This balance makes it well-suited for practical use, ensuring robust and reliable face recognition performance at scale, without imposing excessive computational burdens on typical teacher or institutional hardware.
VI-C A Case Study of iExam’s Post-exam Analysis Report
To illustrate the practical utility of iExam ’s post-exam analysis module, we present a representative case study based on a five-minute exam video. The system automatically generates a set of visual and histogram-style reports, allowing instructors to review exam-taker attendance and identify abnormal behaviors at a glance.
Recognition Frequency Overview. As shown in Fig. 11, the report includes a bar chart summarizing the total number of times each student’s face was correctly recognized during the analysis window. Although the recognition times is computed at the frame level, to improve the testing efficiency, as detailed in Sec. V-B, we treat one successful recognition per student per second as a valid presence record. Since one second typically corresponds to 30 frames, this approach significantly reduces computation while still providing a reliable measure of student presence. The visualization offers a high-level summary of engagement and can help quickly identify students whose faces were rarely detected, potentially flagging technical issues or suspicious behavior.
Figure 11: Conclusion of Recognition Times. The number of times each student’s face was detected and identified over the test interval. To protect privacy, the real names were replaced with generic labels.
Temporal Presence and Absence Analysis. In addition to overall recognition frequency, the system generates fine-grained temporal plots, such as those in Fig. 12, which display the second-by-second presence status for each participant throughout the exam. Each bar segment corresponds to one second where a student was present or absent, allowing instructors to pinpoint the exact times when a student disappeared from view. This feature is particularly useful for cross-checking suspected incidents of absence or face swapping, providing an objective and auditable record for subsequent review.
Figure 12: Summary of recognition distribution. The visualization presents a timeline showing each student’s presence or absence at every second throughout the exam period. To safeguard privacy, real names were anonymized using generic labels.
Consecutive Absence Detection. Finally, as summarized in Fig. 13, the system aggregates the total duration of consecutive absences for each student, highlighting cases where a participant was missing for an extended period. Longer blocks of continuous absence are considered more suspicious and may warrant targeted investigation or follow-up with the corresponding student. This information is presented both as a summary bar chart and can be cross-referenced with the original video for validation.
Figure 13: Consecutive disappear time for testing video. Each bar indicates the longest continuous absence interval detected for each student. For privacy protection, actual names were substituted with generic labels.
By providing automated, interpretable, and time-aligned visual feedback, iExam’s post-exam analysis report enables instructors to efficiently verify student presence, investigate anomalies, and maintain exam integrity with minimal manual effort. The reporting system is scalable to longer videos and larger cohorts, and can be customized or extended for more detailed event logging or alerting as needed.
VII Related Work
Traditional Face Recognition Approaches. Early face recognition research primarily focused on holistic and feature-based statistical methods. The Eigenface approach, based on Principal Component Analysis (PCA) [turk1991eigenfaces, ramadhani2017human], projects face images onto a lower-dimensional subspace spanned by principal components, capturing the main variation across training images. Fisherface [chan2008face] applies Linear Discriminant Analysis (LDA) to maximize between-class separability while minimizing within-class variance, providing robustness against lighting and expression changes. Another commonly used traditional method, Local Binary Patterns Histograms (LBPH) [LBPH], encodes local texture information by thresholding pixel neighborhoods, then aggregates histograms for recognition. Despite their historical significance and efficiency, these methods suffer from degraded performance under varying lighting, pose, and occlusion conditions, and lack the capacity to model non-linear correlations in high-dimensional face data.
Haar-like Features and Classical Detection. Viola and Jones [viola2001rapid] introduced a cascade classifier based on Haar-like features and Adaboost, which became a widely used standard for real-time face detection. Haar-like features capture simple intensity differences between adjacent rectangular regions (e.g., eyes vs. cheeks), leveraging the biological symmetry of faces for efficient object localization. Such approaches, while fast, are limited in their expressiveness and tend to be brittle in unconstrained environments [chen2015enhancing].
OpenCV Face Recognizers. The OpenCV library provides classical, well-established algorithms for face recognition, including the Eigenface, Fisherface, and LBPH methods [Eigenface, Fisherface, LBPH]. These approaches reply primarily on traditional feature extraction and dimensionality reduction techniques rather than deep neural networks. While these implementations are efficient and computationally lightweight, which makes them suitable for small-scale or embedded applications, their performance degrades in complex, real-world scenarios. Online proctoring scenarios introduce challenges such as large pose variations, inconsistent illumination, motion blur, partial occlusion, and low-resolution webcam feeds, all of which severely affect the discriminative power of classical methods. Moreover, these algorithms depend heavily on the quality of face alignment and are limited in their ability to generalize across individuals or sessions without extensive manual calibration. These limitations have motivated the shift toward deep learning-based face recognition models, which often require more computational resources but deliver significantly improved accuracy and robustness. To address these trade-offs, our system incorporates automated model selection, enabling dynamic switching between the classical algorithm, i.e., Haar-cascade method, and Deep Convolutional Neural Networks (DCNNs) based on runtime resource availability and performance requirements.
Deep Learning and CNN-based Approaches. Recent advances in deep learning have revolutionized the field of face recognition and detection. Deep Convolutional Neural Networks (DCNNs) [lecun1998gradient, krizhevsky2012imagenet, he2016deep] learn hierarchical feature representations directly from raw images, automatically capturing complex, non-linear facial characteristics such as pose, illumination, and expression variations. Landmark architectures such as AlexNet [krizhevsky2012imagenet], GoogleNet [szegedy2015going], and ResNet [he2016deep] have demonstrated that increasing network depth and introducing residual connections significantly improve generalization and robustness across large-scale visual recognition benchmarks. Building on these foundations, specialized face recognition models, such as FaceNet [schroff2015facenet] and ArcFace [deng2019arcface], achieve state-of-the-art accuracy by learning discriminative embeddings that map facial images into compact feature spaces. In the context of online proctoring, DCNN-based models provide the accuracy, adaptability, and robustness needed to handle diverse webcam feeds and dynamic exam environments.
Face Detection Algorithms. Several modern deep learning-based detection algorithms have been developed, including Multi-task Cascaded Convolutional Networks (MTCNNs) [zhang2016joint], Single Shot-Multibox Detector (SSD) [liu2016ssd], and DCNN models integrated with frameworks such as OpenCV. These algorithms have proven highly effective in challenging real-world conditions and they can accurately localize faces in complex scenes and are less sensitive to pose and illumination changes compared to classical methods. Our system evaluates and leverages the performance and efficiency trade-offs of these mainstream detectors for both real-time and post-exam analysis.
Optical Character Recognition (OCR) in Human Identity Verification. OCR plays a pivotal role in scenarios where automatic identity extraction from video is needed. State-of-the-art OCR engines, such as Google’s Tesseract (used via PyTesseract in Python), perform multi-stage processing including image acquisition, denoising, binarization, segmentation, feature extraction, and text recognition. They are widely adopted in academic and industrial applications for extracting textual identity cues from screen overlays or exam platforms [tesseract_github]. Recent research also explores integrating OCR with face recognition to enable automatic ground-truth labeling in large-scale datasets [arlazarov2018midv, bulatov2021midv], a technique we adopt to facilitate hands-free, dynamic identity annotation in online exams.
Automated Proctoring Systems. A number of recent works and commercial solutions (e.g., ProctorU and open-source frameworks) aim to automate the monitoring of remote exams via video analysis, using face detection, face recognition, gaze estimation, and anomaly detection [proctoru_official_website, butler2020systematic]. However, few systems provide fine-grained, two-stage separation of lightweight real-time detection and deep post-exam analysis as in iExam. Our method uniquely addresses the trade-off between speed, privacy, and post-exam integrity verification by combining client-side lightweight detection with deep learning-powered offline recognition and reporting.
In summary, while classical statistical and feature-based methods laid the groundwork for face analysis, the state-of-the-art has shifted toward deep learning-based models for robust performance under realistic, unconstrained conditions. Our approach leverages this paradigm shift, integrating efficient DCNN-based detection and recognition, automated OCR labeling, and tailored analysis for online exam proctoring, to deliver a not only accurate and scalable but also privacy-preserving solution.
VIII Conclusion
This paper presented iExam, an effective online exam proctoring system that combines lightweight real-time face detection with deep learning-based post-exam face recognition. By decoupling instant monitoring from in-depth analysis, iExam achieves both efficiency and high accuracy while protecting user privacy. Experimental results demonstrate that the system delivers robust performance on real-world exam videos, enabling reliable detection of student presence and identification of suspicious behaviors. Our findings suggest that iExam offers a practical and scalable solution for secure and transparent online exam supervision.
Acknowledgments
This research/project was partially supported by UGC-funded Courseware Development Grant (ref. no. 4170827) and Funding Scheme for Virtual Teaching and Learning (ref. no. 4170890) from The Chinese University of Hong Kong. This research has also benefited from financial support of Lingnan University, Hong Kong Special Administrative Region, China.